OpenStudy (sedatefrog712):

Medal The length of a rectangle is (4x2 + 2x - 1) units, and its width is (3x3 - x + 4) units. Part A: What is the area of the rectangle? Show your work Part B: Does the answer for Part A show that polynomials are closed under an operation? Justify your answer Part C: What is the degree and classification of the expression obtained in Part A?

OpenStudy (sedatefrog712):

@triciaal

OpenStudy (sedatefrog712):

would I use like terms

OpenStudy (triciaal):

the area of a rectangle is length * width A = L* W

OpenStudy (sedatefrog712):

o ok so I would use the distributive property

OpenStudy (triciaal):

you will multiply the length by the width. you use like terms to simplify calculations

OpenStudy (sedatefrog712):

12x^5-2x-4

OpenStudy (triciaal):

yes use distributive property multiply each term inside by what is outside

OpenStudy (triciaal):

you did that too quick

OpenStudy (triciaal):

(4x2 + 2x - 1)(3x3 - x + 4) =?

OpenStudy (triciaal):

4 x ^2(3x3 - x + 4) + 2x(3x3 - x + 4) -1(3x3 - x + 4) = ?

OpenStudy (triciaal):

is this what you did?

OpenStudy (sedatefrog712):

12x^5 -4x + 16x^2 6x^4 -2 + 8x -3x^3 + x -4

OpenStudy (sedatefrog712):

this is for 4x^2 2x -1

OpenStudy (sedatefrog712):

then i add them together?

OpenStudy (triciaal):

you have 12x^5 -4x + 16x^2 that should not be -4x

OpenStudy (triciaal):

you have a problem getting the middle term when you multiply by an x outside the parenthesis

OpenStudy (triciaal):

4x^2 * (-x) = -4x^3 and 2 x*(-x) = -2 x^2

OpenStudy (sedatefrog712):

o ok

OpenStudy (sedatefrog712):

12x^5 -4x^3 + 16x^2 6x^4 -2x@ + 8x -3x^3 + x -4

OpenStudy (triciaal):

check your typo and this is all one expression so group like terms to simplify (when you submit it's all on one line)

OpenStudy (sedatefrog712):

ok thank you I gtg now bye

OpenStudy (triciaal):

ok